Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis ofReq 1 Req 2 Req 3A Req 3B Req 4 Prepare a standard cost card for the companys product. (Round your answers to 2 decimal placReq 1 Req 2 Req 3A Req 3B Req 4 Compute the standard direct labor-hours allowed for the years production. Standard direct laReq 1 Req 2 Req 3A Req 3B Req 4 Complete the following Manufacturing Overhead T-account for the year. Manufacturing OverheadReq 1 Req 2 Req 3A Req 3B Req 4 Determine the reason for any underapplied or overapplied overhead for the year by computing t
Answer #1

Solution 1:

Computation of Predetermined overhead rate - Lane Company
Particulars Amount
Budgeted variable manufacturing overhead ($5.40*285000) $15,39,000
Budgeted fixed manufacturing overhead $26,79,000
Total Manufacturing overhead $42,18,000
/Budgeted direct labor hours 285000
Predetermined overhead rate $14.80
Less: Variable overhead rate (Per direct labor hour) $5.40
Fixed manufacturing overhead rate (per direct labor hour) $9.40

Solution 2:

Standard Cost Card - Lane Company
Direct Material 4 Pounds at $11.50 per pound $46.00
Direct labor 1.5 DLHs at $13.70 per DLH $20.55
Variable manufacturing overhead 1.5 DLHs at $5.40 per DLH $8.10
Fixed manufacturing overhead 1.5 DLHs at $9.40 per DLH $14.10
Standard cost per unit $88.75

Solution 3a:

Standard hours allowed = 228000*1.5 = 342,000 hours

Solution 3b:

Manufacturing Overhead
Particulars Debit Credit Particulars
Actual costs (1148550+ 2964000) $41,12,550 $50,61,600 Applied costs (342000*$14.80)
Overapplied overhead $9,49,050
Total $50,61,600 $50,61,600 Total

Solution 4:

Variable overhead actual rate = $1148550/ 370500 = $3.10 per hour
Variable overhead rate variance = (SP - AP) *Actual Hours = ($5.40 - $3.10) * 370500 = $852,150 Favorable
Variable overhead Efficiency variance = (Standard hours - Actual Hours) *SP = (342000 - 370500)*$5.40 = $153,900 (Unfavorable)
Fixed overhead budget Variance = Budgeted Fixed Overhead - Actual Fixed overhead = $2679000 - $2964000 = $285,000 (unfavorable)
Fixed Overhead Volume Variance = ($9.40*342000) - $2679000 = $535,800 Favorable